tennis player made a successful first serve 67% of the times. If she continues to serve at the same rate the next time she play
and serves 6 times in the first game, determine the probability that:
1.2.1. All six first serves will be in.
1.2.2. At least 4 first serves will be in.
1
Expert's answer
2020-03-06T15:17:25-0500

It is a Binomial distribution problem with "P=0.67, n=6"

1.2.1 "P(=6)"

The Excel Binomial formula=BINOM.DIST(6,6,0.67,FALSE) yields 0.090458.

Answer: 0.090458

1.2.2 "P(\geq4)"

"P(\ge4)=1-P(\le3)"

From BINOM.DIST Excel formula; "P(\le3)" is given by =BINOM.DIST(3,6,0.67,TRUE) which yields 0.313048

Thus, "P(\ge4)=1-0.313048=0.686952"